# Monotonicity of Recurrence in Random Walks

### v2

@@ Abstract.
We consider non-homogeneous random walks on the positive quadrant in two dimensions. In the 1960’s the following question was asked: is it true if such a random walk $X$ is recurrent and $Y$ is another random walk that at every point is more likely to go down and more likely to go left than $Y$, then $Y$ is also recurrent?
We provide an example showing that the answer is negative. We also show that if either the random walk $X$ or $Y$ is sufficiently homogeneous then the answer is in fact positive.
@@ 1. Introduction
For the non-homogeneous reflecting random walk on $\mathbb{N}$ with transitions $p_{n}$ from $n$ to $n+1$ and $q_{n}=1-p_{n}$ from $n$ to $n-1$ (for $n>0$) there is an explicit characterization of when the random walk is recurrent/transient (see [2, Section XV.8]). From this it follows directly that if the random walk is recurrent and we consider a different one in which for all $n$ the $p_{n}^{\prime}$ satisfy $p_{n}^{\prime}\leq p_{n}$ then the new random walk is also recurrent, while if the first was transient while the second satisfies $p_{n}^{\prime}\geq p_{n}$ for all $n$ then the second is also transient. These monotonicity properties can also be easily seen via a coupling argument.
In the early 1960’s the question as to whether or not a similar monotonicity property holds for two dimensional random walks circulated among the probabilists at Princeton University. To the best of our knowledge this question has not been resolved to date. The purpose of the following note is to resolve this question.
@@
For two random walks $X$ and $Y$ on $\mathbb{N}^{2}$, say $X\preceq Y$ if for each state, the probability of going down is weakly greater in $X$ than in $Y$, the probability of going left is weakly greater in $X$ than in $Y$, the probability of going up is weakly greater in $Y$ than in $X$, and the probability of going right is weakly greater in $Y$ than in $X$.
@@
Note that for two irreducible Markov chains $X$ and $Y$ on the same state space, if there only exist finitely many states for which the outgoing transition probabilities are different for $X$ and $Y$, i.e., $X$ and $Y$ agree on a cofinite set of states, then they will have the same recurrence behavior: $X$ is recurrent, positive recurrent, or transient if and only if $Y$ is recurrent, positive recurrent, or transient, respectively. Thus, when considering our question of monotonicity of recurrence, we can generalize Definition 1.1, as well as all other definitions that follow in this paper, so that the required inequalities only need to hold on a cofinite set of states. However, for sake of clarity and brevity, we henceforth do not discuss this equivalence up to a cofinite set.
The intuition behind this question is that if $Y$ is recurrent then it returns to the origin with probability $1$. As $X\preceq Y$ it is perhaps natural to expect that $X$ will also return to the origin with probability $1$, as indeed is the case in dimension one.
@@
On the other hand, we find conditions under which recurrence is monotone.
@@
$\begin{cases}\text{right}&\text{w.p. }r_{o}\\ \text{up}&\text{w.p. }u_{o},\end{cases}\qquad\begin{cases}\text{left}&\text{w.p. }\ell_{x}\\ \text{right}&\text{w.p. }r_{x}\\ \text{up}&\text{w.p. }u_{x},\end{cases}\qquad\begin{cases}\text{down}&\text{w.p. }d_{y}\\ \text{up}&\text{w.p. }u_{y}\\ \text{right}&\text{w.p. }r_{y},\end{cases}\qquad\begin{cases}\text{left}&\text{w.p. }\ell_{q}\\ \text{right}&\text{w.p. }r_{q}\\ \text{up}&\text{w.p. }u_{q}\\ \text{down}&\text{w.p. }d_{q},\end{cases}$
@@
$r_{o}\geq r_{x}=r_{q},\quad\ell_{x}\geq\ell_{q},\quad u_{x}\geq u_{q},\quad u_{o}\geq u_{y}=u_{q},\quad d_{y}\geq d_{q},\quad\text{and}\quad r_{y}\geq r_{q}.$
@@
We now show that under an inward-homogeneity condition, recurrence and positive recurrence are monotonic properties with respect to $\preceq$.
@@
If an irreducible inward-homogeneous random walk $X$ on $\mathbb{N}^{2}$ is recurrent (respectively, positive recurrent), then any irreducible random walk $Y$ on $\mathbb{N}^{2}$ such that $Y\preceq X$ is also recurrent (respectively, positive recurrent).
For the reverse direction, where we want transience of $X$ to imply transience of $Y$ for $Y\succeq X$, we can of course take the contrapositive of Theorem 1.5, but this requires $Y$ to be inward-homogeneous. The following theorem instead assumes $X$ is inward-homogeneous, and its proof is very similar to that of Theorem 1.5.
@@
If an irreducible inward-homogeneous random walk $X$ on $\mathbb{N}^{2}$ is transient, then any irreducible random walk $Y$ on $\mathbb{N}^{2}$ such that $Y\succeq X$ is also transient.
Using Theorems 1.5 and 1.6 and their contrapositives, we obtain the following result implying monotonicity when neither random walk is inward-homogeneous, though an inward-homogeneous random walk must exist between the two random walks.
@@
Let $X$, $Y$, and $Z$ be irreducible random walks on $\mathbb{N}^{2}$ where $Y$ is inward-homogeneous and $X\preceq Y\preceq Z$. If $Z$ is recurrent then $X$ is recurrent, and if $X$ is transient then $Z$ is transient.
@@
A similar monotonicity question can be asked for slabs, i.e., sets of the form $\{(i,j)\in\mathbb{Z}^{2}:i\geq 0,0\leq j\leq k\}$, which can be viewed as an interpolation between the one-dimensional environment of non-homogeneous random walks on $\mathbb{N}$, for which monotonicity always holds trivially, and the two-dimensional environment that we study. We obtain analogous results for slabs in Section 4.
@@
Rayleigh’s monotonicity law, introduced by Rayleigh [3] in 1899, may be able to make some statements on this question of when monotonicity of recurrence occurs, but is severely limited by the requirement that both $X$ and $Y$ must be Markov chains that correspond to electric networks (see [1] for a comprehensive overview). A Markov chain corresponds to an electric network if there exist constants $c_{xy}\geq 0$ for every unordered pair of two distinct states $x$ and $y$, such that letting $c_{x}=\sum_{y}c_{xy}$ for all states $x$, assuming only finitely many terms in each sum are positive, we have the transition probability from $x$ to $y$ is $c_{xy}/c_{x}$ and the transition probability from $y$ to $x$ is $c_{xy}/c_{y}$. When the Markov chain is positive recurrent, this is equivalent to the Markov chain being reversible, where the stationary distribution $\pi$ is given by $\pi_{x}\propto c_{x}$. This imposes certain conditions on a random walk on $\mathbb{N}^{2}$: for example, for 4 states $w$, $x$, $y$, and $z$, we have $\frac{c_{wx}}{c_{xy}}\cdot\frac{c_{xy}}{c_{yz}}\cdot\frac{c_{yz}}{c_{wz}}\cdot\frac{c_{wz}}{c_{wx}}=1$. Letting $(w,x,y,z)=((0,0),(1,0),(1,1),(0,1))$ in an inward-homogeneous random walk on $\mathbb{N}^{2}$, this implies $\frac{\ell_{x}d_{q}r_{y}u_{o}}{u_{x}\ell_{q}d_{y}r_{o}}=1$. Considering two other “squares” $((1,0),(2,0),(2,1),(1,1))$ and $((0,1),(1,1),(1,2),(0,2))$ yields $\frac{r_{x}\ell_{q}}{\ell_{x}r_{q}}=1$ and $\frac{u_{q}d_{y}}{u_{y}d_{q}}=1$. These conditions severely limit the applicability of Rayleigh monotonicity to two random walks $X\preceq Y$.
In Section 2, we prove Theorem 1.3. In Section 3, we prove Theorems 1.5 and 1.6. Lastly, in Section 4, we prove analogues of Theorems 1.3, 1.5, 1.6 and 1.7 for slabs.

### v4

# Monotonicity, Topology, and Convexity of Recurrence in Random Walks

@@ Abstract.
We consider non-homogeneous random walks on the two-dimensional positive quadrant $\mathbb{N}^{2}$ and the one-dimensional slab $\{0,1,\dots,k\}\times\mathbb{N}$. In the 1960’s the following question was asked for $\mathbb{N}^{2}$: is it true if such a random walk $X$ is recurrent and $Y$ is another random walk that at every point is more likely to go down and more likely to go left than $Y$, then $Y$ is also recurrent?
We provide an example showing that the answer is negative. We also show, via a coupling argument, that if either the random walk $X$ or $Y$ is sufficiently homogeneous then the answer is in fact positive. In addition, we show using the Rayleigh monotonicity principle that the analogous question for random walks on trees is positive.
These results show that the subset of parameter space that yields recurrent random walks possesses some geometric properties, in this case the structure of an order ideal. Motivated by this perspective, we consider the more symmetric setting of homogeneous random walks on finitely generated abelian groups, and ask when this subset possesses other geometric properties, namely various topological properties and convexity. We answer some of these questions: in particular, we show that this subset is closed, and under a symmetric support condition, show it is path-connected and additionally show it is convex if and only if its effective dimension is at most 2. We also show its complement is in some sense typically path-connected but not convex. We finally propose some related open problems.
@@ 1. Introduction
For the non-homogeneous reflecting random walk on $\mathbb{N}$ with transitions $p_{n}$ from $n$ to $n+1$ and $q_{n}=1-p_{n}$ from $n$ to n1n-1 (for $n>0$) there is an explicit characterization of when the random walk is recurrent/transient (see [11, Section XV.8]). From this it follows directly that if the random walk is recurrent and we consider a different one in which for all $n$ the $p_{n}^{\prime}$ satisfy $p_{n}^{\prime}\leq p_{n}$ then the new random walk is also recurrent, while if the first was transient while the second satisfies $p_{n}^{\prime}\geq p_{n}$ for all $n$ then the second is also transient. These monotonicity properties can also be easily seen via a coupling argument.
In the early 1960’s the question as to whether or not a similar monotonicity property holds for two-dimensional random walks circulated among the probability group at Princeton University. We are not sure if this is the first time this question was asked. To the best of our knowledge this question has not been resolved to date, though it is not hard to derive the answer using a 1972 result by Malyshev [16]; see Remark 2.1. The purpose of the following note is to resolve this question and related questions.
@@
For two random walks $X$ and $Y$ on $\mathbb{N}^{2}$, say $X\preceq Y$ if for each state, the probability of going down in $X$ is greater than or equal to that in $Y$, the probability of going left in $X$ is greater than or equal to that in $Y$, the probability of going up in $Y$ is greater than or equal to that in $X$, and the probability of going right in $Y$ is greater than or equal to that in $X$.
@@
Note that for two irreducible Markov chains $X$ and $Y$ on the same state space, if there only exist finitely many states for which the outgoing transition probabilities are different for $X$ and $Y$, i.e., $X$ and $Y$ agree on a co-finite set of states, then they will have the same recurrence behavior: $X$ is recurrent, positive recurrent, or transient if and only if $Y$ is recurrent, positive recurrent, or transient, respectively. Thus, when considering our question of monotonicity of recurrence, we can generalize Definition 1.1, as well as all other definitions that follow in this paper, so that the required inequalities only need to hold on a co-finite set of states. For example, we can ignore any inequalities concerning the transition probabilities from the origin. However, for sake of clarity and brevity, we henceforth do not discuss this equivalence up to a co-finite set.
The intuition behind 1.2 is that if $Y$ is recurrent then it returns to the origin with probability $1$. As $X\preceq Y$ it is natural to expect that $X$ will also return to the origin with probability $1$, as indeed is the case in nearest-neighbor walks on $\mathbb{N}$ as previously discussed.
@@
Our construction (see Figure 1 for a schematic) is of a similar nature as the following perhaps more intuitive example of a non-nearest-neighbor walk on $\mathbb{N}$, for which we only sketch the main idea. Suppose we have a random walk $Y$ on $\mathbb{N}$ whose jumps take values in $\left\{-2,-1,1,2\right\}$, and whose odd states drift towards 0 and even states drift towards $\infty$, in such a way that the walk tends to stay on odd states enough that $Y$ is positive recurrent. Then, one can make each state drift slightly more towards the left but in such a way that the walk tends to stay on the even states, so that the new Markov chain $X$ is transient. Our construction essentially uses the positive $x$-axis as the escape route, analogous to the even states, while the positive $y$-axis plays the role of the odd states.
On the other hand, we find conditions under which recurrence is monotone with respect to $\preceq$.
@@
[…]
@@
(1.2) $r_{o}\geq r_{x}=r_{q},\quad\ell_{x}\geq\ell_{q},\quad u_{x}\geq u_{q},\quad u_{o}\geq u_{y}=u_{q},\quad d_{y}\geq d_{q},\quad\text{and}\quad r_{y}\geq r_{q}.$
A random walk on $\mathbb{N}^{2}$ of the above form (1.1), though not necessarily satisfying (1.2), is commonly referred to as maximally homogeneous in the literature, and is the one most studied in the literature for random walks on the quadrant; see, e.g., [7, 8, 10]. Typically, maximal homogeneity is defined to only require the mean drifts to be the same in the four regions as opposed to requiring the transition probabilities to be the same.
@@
Remark 1.5.
Comparison ideas are widely used in the queuing literature; see, e.g., [22] or its English translation [23]. The queuing setup, specifically the Jackson network of two M/M/1 queues, translates to a continuous time random walk on $\mathbb{N}^{2}$, where movement in this continuous time random walk is dictated by Poisson rates, and boundary conditions are enforced by simply suppressing illegal moves, i.e., waiting until the Poisson clock of a valid move rings. Motivated by this option of being lazy instead of redirecting the illegal move, we introduce the following weaker condition in Definition 1.6. Essentially, instead of applying redirection, e.g., redirecting from down to the left or up when on the positive $x$-axis, one can choose to not move, which in our non-lazy random walk setting corresponds to reallocating some of the down probability $d_{q}$ (in the case of the positive $x$-axis) to the other three directions proportional to $\ell_{q}$, $u_{q}$, and $r_{q}$. Because the following definition allows for both suppressing or redirecting the illegal move, we found that introducing the intermediate definition of inward-homogeneity was still instructive for intuition.
Definition 1.6.
Using the notation of Definition 1.4, i.e., (1.1), say a random walk on $\mathbb{N}^{2}$ is weakly inward-homogeneous if
(1.3) $\frac{\ell_{x}}{\ell_{q}},\frac{u_{x}}{u_{q}}\geq\frac{r_{x}}{r_{q}}\geq 1\quad\text{and}\quad\frac{d_{y}}{d_{q}},\frac{r_{y}}{r_{q}}\geq\frac{u_{y}}{u_{q}}\geq 1.$
In our positive result we show that under a weakly inward-homogeneity condition, recurrence and positive recurrence are monotonic properties with respect to $\preceq$.
@@
If an irreducible weakly inward-homogeneous random walk $X$ on $\mathbb{N}^{2}$ is recurrent (respectively, positive recurrent), then any irreducible random walk $Y$ on $\mathbb{N}^{2}$ such that $Y\preceq X$ is also recurrent (respectively, positive recurrent).
For the reverse direction, where we want transience of $X$ to imply transience of $Y$ for $Y\succeq X$, we can of course take the contrapositive of Theorem 1.7, but this requires $Y$ to be weakly inward-homogeneous. The following theorem instead assumes $X$ is weakly inward-homogeneous, and its proof is very similar to that of Theorem 1.7.
@@
If an irreducible weakly inward-homogeneous random walk $X$ on $\mathbb{N}^{2}$ is transient, then any irreducible random walk $Y$ on $\mathbb{N}^{2}$ such that $Y\succeq X$ is also transient.
Using Theorems 1.7 and 1.8 and their contrapositive, we obtain the following result implying monotonicity when neither random walk is weakly inward-homogeneous, though a weakly inward-homogeneous random walk must exist between the two random walks.
@@
Let $X$, $Y$, and $Z$ be irreducible random walks on $\mathbb{N}^{2}$ where $Y$ is weakly inward-homogeneous and $X\preceq Y\preceq Z$. If $Z$ is recurrent then $X$ is recurrent, and if $X$ is transient then $Z$ is transient.
Remark 1.10.
The classification of recurrence and transience for random walks in the quadrant $\mathbb{N}^{2}$ goes back to Kingman [12] in 1961 and Malyshev [14, 16] in the early 1970s. As discussed in further detail in Remark 2.1, this work is able to immediately imply the two walks constructed for the proof of Theorem 1.3 are respectively positive recurrent and transient, though this work did not comment on the non-monotonicity of recurrence, e.g., by providing an example demonstrating monotonicity fails. The Kingman–Malyshev result provides an essentially complete classification in the maximally homogeneous case where the interior drift is non-zero, with the classification depending on the drift vector in the interior and the two angles produced by the drifts at the boundaries. Malyshev’s hypothesis of bounded increments was relaxed through many works [9, 19, 8, 24, 26] to only require finite first moments. The essentially complete recurrence classification in the case of zero interior drift took much longer, until the early 1990s [2, 7, 8], and is more complex, depending on the increment covariance matrix in the interior as well as the two boundary reflection angles.
The recurrence classification in the non-zero interior drift case is simple enough that it should essentially imply the monotonicity result of Theorem 1.7 in this case; one can obtain a quadratic Lyapunov function for showing positive recurrence, where the level curves are like those in [8, Figure 3.3.1]. The zero interior drift case is more delicate, as we must consider the interior covariance, but more recent work, e.g., [4], may be able to deduce a similar result by applying a suitable transformation to put the walk into a wedge. An advantage of these Foster–Lyapunov methods is that they depend only on the drifts (and sometimes the covariances), providing the potential to expand our results beyond nearest-neighbor walks. In the nearest-neighbor case, however, our coupling approach to prove Theorems 1.7 and 1.8 is simple and addresses both cases simultaneously.
We hope our scheme of comparing non-homogeneous random walks, where the general case seems entirely intractable, to tractable cases, such as weakly inward-homogeneous walks in Corollary 1.9, is a fruitful idea for further work. Our paper establishes this comparison via a coupling approach, but these tractable cases, e.g., maximally homogeneous walks, are where Foster–Lyapunov methods are also successful. We welcome further work on this topic and leave the application of different methods to our comparison scheme as an open problem for an interested reader.
Remark 1.11.
We believe that with additional work, one may be able to extract more from our coupling idea in the positive recurrent case, i.e., the proof of Theorem 1.7 in Section 3, namely obtaining comparisons of tail bounds for return times for the two walks, e.g., inequalities of the form $\mathbb{P}(\tau\geq n)\leq C\mathbb{P}(\tau^{\prime}\geq an+b)$ for constants $C$, $a$, and $b$. We leave this as an open problem. Such tail bounds help compute moments of passage times, which are important for numerous reasons, including mixing rates, tails of stationary measures, and more. In the case of the quadrant, bounds on moments of passage times in the maximally homogeneous case are given by Aspandiiarov, Iasnogorodski, and Menshikov [1], and we refer readers to [4] for a more recent account. However, these methods are unable to directly compare the tail probabilities of two such walks, and so establishing such comparison inequalities, perhaps through a more detailed analysis of this coupling, would be of interest.
In the special case of Theorem 1.7 where $X$ has $u_{x}=u_{q}$ and $r_{y}=r_{q}$, i.e., $X$ is inward-homogeneous and its redirection rules also forbid the positive $x$-axis redirecting up and similarly the positive $y$-axis redirecting right, so that they can only redirect left or down, respectively, if $\tau_{X}$ and $\tau_{Y}$ denote the return times of the origin in $X$ and $Y$, respectively, it is straightforward to modify our proof of Theorem 1.7 to see that we can couple $X$ and $Y$ such that $Y_{t}\leq X_{t}$ for all $t$ and thus $\tau_{Y}\leq\tau_{X}$, yielding the comparison inequality $\mathbb{P}(\tau_{Y}\geq n)\leq\mathbb{P}(\tau_{X}\geq n)$ for all $n$.
@@
A similar monotonicity question can be asked for slabs $\{(i,j)\in\mathbb{Z}^{2}:i\geq 0,0\leq j\leq k\}$, which can be viewed as an interpolation between the one-dimensional environment of non-homogeneous random walks on $\mathbb{N}$, for which monotonicity always holds trivially, and the two-dimensional environment that we study. We obtain analogous results for slabs in Section 4.
Remark 1.12.
@@
Rayleigh’s monotonicity law, introduced by Rayleigh [18] in 1899, may be able to make some statements on this question of when monotonicity of recurrence occurs, but is severely limited by the requirement that both $X$ and $Y$ must be Markov chains that correspond to electric networks (see [6] for a comprehensive overview). A Markov chain corresponds to an electric network if there exist constants $c_{xy}\geq 0$ for every unordered pair of two distinct states $x$ and $y$, such that letting $c_{x}=\sum_{y}c_{xy}$ for all states $x$, assuming only finitely many terms in each sum are positive, we have the transition probability from $x$ to $y$ is $c_{xy}/c_{x}$ and the transition probability from $y$ to $x$ is $c_{xy}/c_{y}$. When the Markov chain is positive recurrent, this is equivalent to the Markov chain being reversible, where the stationary distribution $\pi$ is given by $\pi_{x}\propto c_{x}$. This imposes certain conditions on a random walk on $\mathbb{N}^{2}$: for example, for 4 states $w$, $x$, $y$, and $z$, we have $\frac{c_{wx}}{c_{xy}}\cdot\frac{c_{xy}}{c_{yz}}\cdot\frac{c_{yz}}{c_{wz}}\cdot\frac{c_{wz}}{c_{wx}}=1$. Letting $(w,x,y,z)=((0,0),(1,0),(1,1),(0,1))$ in an inward-homogeneous random walk on $\mathbb{N}^{2}$, this implies $\frac{\ell_{x}d_{q}r_{y}u_{o}}{u_{x}\ell_{q}d_{y}r_{o}}=1$. Considering two other “squares” $((1,0),(2,0),(2,1),(1,1))$ and $((0,1),(1,1),(1,2),(0,2))$ yields $\frac{r_{x}\ell_{q}}{\ell_{x}r_{q}}=1$ and $\frac{u_{q}d_{y}}{u_{y}d_{q}}=1$. These conditions severely limit the applicability of Rayleigh monotonicity to two random walks $X\preceq Y$. The restrictiveness of the setting of Rayleigh monotonicity, i.e., reversibility, can also easily be seen in the case of non-nearest-neighbor walks on $\mathbb{N}$.
However, one can use Rayleigh monotonicity to show monotonicity of recurrence holds for trees, in the following sense. First, for two random walks $X$ and $Y$ on the same (undirected) tree rooted at $r$, there are two natural ways to define $X\preceq Y$ analogous to our definition for random walks on $\mathbb{N}^{2}$: we can either simply require that for all states $v\neq r$, the probability of transitioning from $v$ to its parent in $X$ is greater than or equal to that in $Y$; or, we can additionally require that for each state $v$ and for any of its children $w$, the transition probability of $v$ to $w$ in $X$ is less than or equal to that in $Y$. Under the former definition, which does not control the relative balance of the transition probabilities from a state to its children, one can construct a counterexample to monotonicity of recurrence, i.e., an example $X\preceq Y$ such that $X$ is transient yet $Y$ is (positive) recurrent. We provide such a counterexample in Theorem 5.3. However, under the latter definition, one can use Rayleigh monotonicity to show that monotonicity of recurrence—as well as monotonicity of the expected return time to the root for positive recurrent chains—holds in general; we establish this in Theorem 5.4.
These results on $\mathbb{N}^{2}$, the slab, and trees show that the subset of parameter space that yields recurrent random walks possesses some geometric properties, in this case the structure of an order ideal with respect to some natural partial order. Motivated by this perspective, we consider the more symmetric setting of homogeneous random walks on finitely generated abelian groups. For a finitely generated abelian group $G$, we can identify the homogeneous random walks that transition according to some probability distribution over some fixed finite subset $S\subseteq G$ with the probability simplex on $S$. We then consider the subset $R$ of this simplex that corresponds to recurrent random walks, and ask when it possesses certain geometric properties, namely various topological properties and convexity. We answer some of these questions in Section 6: in particular, we show that $R$ is closed, and if $S$ is symmetric, show $R$ is path-connected and additionally show $R$ is convex if and only if its effective dimension is at most 2 (see Corollary 6.6 for the precise formulation). We also show its complement $R^{c}$ is in some sense typically path-connected but not convex.
Remark 1.13.
We comment on some of the similarities between our work and that on homogeneous random walks on $\mathbb{R}^{d}$, i.e., sums of i.i.d. random variables with distribution $\mu$. In this setting, the Chung–Fuchs theorem [3] completely classifies when the random walk is recurrent, but questions concerning monotonicity and convexity are nontrivial in general. Of course, when $d\geq 3$ and $\mu$ is nondegenerate with mild regularity, e.g., finite second moments certainly suffice, the Chung–Fuchs theorem implies all walks are transient, while for $d\in\left\{1,2\right\}$, if the $d$th moment exists and $\mu$ has mean 0, then the walk is recurrent, so the situation is clear. In general, however, when moments need not exist, Mineka [17] showed for $d=1$ that mixtures of recurrent random walks need not be recurrent, even when $\mu$ is symmetric, demonstrating an example of non-convexity. And in terms of monotonicity, Shepp [20], later generalized by Dharmadhikari and Joag-dev [5], showed that for unimodal transition laws $F$ and $G$, if $F$ is less peaked than $G$, then recurrence of $F$ implies recurrence of $G$.
In Section 2, we prove Theorem 1.3, and in Section 3, we prove Theorems 1.7 and 1.8. We prove analogues of Theorems 1.3, 1.7, 1.8 and 1.9 for slabs, i.e., state spaces of the form $\{(i,j)\in\mathbb{Z}^{2}:i\geq 0,0\leq j\leq k\}$ for some positive integer $k$, in Section 4. We prove the results about trees, Theorems 5.3 and 5.4, in Section 5. Lastly, in Section 6 we consider whether the subset $R$ of parameter space that yields recurrent random walks, for homogeneous random walks on finitely generated abelian groups, possesses various geometric and topological properties.